Bads is thinking of two numbers .one number is 7more than the other number .the product of the number is 60.what are the numbers?
Answer by checkley79(3341)

X(X+7)=60
X^2+7X-60=0
(X-5)(X+12)=0
X-5=0
X=5 ANS. FOR THE SMALLER NUMBER.
5+7=12 ANS FOR THE LARGER NUMBER.
PROOF:
5*12=60
60=60